OpenStudy (anonymous):

Optimization. Find the minimum distance of a point on the graph xy^2=16 from the origin

OpenStudy (anonymous):

is it (xy)^2 or xy^2 if it is the latter then assume that the coordinates( t, 4/sqrt(t)) .. then d = sqrt ( t^2 + 16/t^2) .... differentiating ... 2t -32/t^3 = t = 2 (2,2*sqrt(2))